### v1

# Benchmarking quantum computers

Abstract
The rapid pace of development in quantum computing technology has sparked a proliferation of benchmarks for assessing the performance of quantum computing hardware and software. Good benchmarks empower scientists, engineers, programmers, and users to understand a computing system’s power, but bad benchmarks can misdirect research and inhibit progress. In this Perspective, we survey the science of quantum computer benchmarking. We discuss the role of benchmarks and benchmarking, and how good benchmarks can drive and measure progress towards the long-term goal of useful quantum computations, i.e., “quantum utility”. We explain how different kinds of benchmark quantify the performance of different parts of a quantum computer, we survey existing benchmarks, critically discuss recent trends in benchmarking, and highlight important open research questions in this field.
In quantum computing, benchmarking means measuring the performance of quantum computing systems, and benchmarks are methods that are used to measure performance (Fig. 1a). The rapid advance of quantum computing hardware over the past decade has spawned a bewildering variety of concepts and techniques within the fledgling science of quantum computer benchmarking. In 2014, cutting edge “quantum computers” were physics experiments on a few qubits [1]. Today’s quantum computers routinely deploy 20-400 qubits [2, 3, 4, 5, 6] and can outperform classical computers on specially designed tasks [2]. But despite this progress, quantum computers are still in their infancy. Our field’s goal of solving problems of practical significance—often called “quantum utility”—remains elusive, and many scientists predict that achieving quantum utility will require years of progress and major technological advances [7, 8, 9, 10]. Benchmarking and benchmarks can measure and guide progress toward that goal.
Figure 1: Quantum computer benchmarks. a. Quantum computer benchmarks are methods that are run on quantum computing systems, or on some of their subsystems (qubits, compilers, etc), to measure performance. Each benchmark measures one or more metrics of performance, such as the error rates of a system’s quantum gates. b. Benchmarks enable comparing a quantum computer’s performance to other contemporary, historical, or hypothetical systems, e.g., hypothetical systems on a particular roadmap to quantum utility.
Benchmarks for classical computers, such as the LINPACK [11] and SPEC [12] benchmarks, measure proxies for computational utility. Once quantum computers become capable of computational utility, many quantum computer benchmarks will too. But today, and in the near term, quantum computer benchmarks should instead measure, guide, and incentivize progress towards utility (Fig. 1b). To solve practical problems that classical computers cannot, it is expected that quantum systems will need to grow to millions of physical qubits [10, 7, 9, 8]. But more qubits is not enough. Achieving quantum utility requires reducing and mitigating hardware errors, which are the imperfections in qubits and quantum logic gates that cause large quantum programs to fail (see Box 1). The importance of other limiting factors, like speed or power consumption, pales in comparison. Quantum logic operations fail far more often (today, typically around 0.01%-1% of the time [2, 3, 4, 5, 6]) than classical instructions (around $10^{-23}\%$ of the time [13]), and correcting or mitigating quantum errors is startlingly hard [14, 15, 16]. Because errors will dominate the performance and design of quantum computers for the foreseeable future, most quantum computer benchmarks measure the impacts of those errors.
Many of the earliest and most widely-used quantum computer benchmarks, such as randomized benchmarking (RB) [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], directly measure the rates of errors in a quantum computer’s quantum logic gates. But quantum computers are not just arrays of qubits. Instead, a phalanx of powerful classical computing subcomponents—controllers, compilers, routers, and schedulers—manage and control the qubits [30]. Increasingly, quantum computations will be performed not directly on physical qubits, but instead on higher-performing logical qubits encoded within many physical qubits running a quantum error correction (QEC) algorithm [14]. These complexities have motivated an ever-increasing array of incomparable and complementary benchmarks that test different subsystems in a quantum computer and/or measure different performance metrics. Understanding and leveraging a benchmark requires knowing what systems it tests, what it measures about those subsystems’ performance, and why that metric of performance matters.
In this perspective, we survey the science of quantum computer benchmarking and we highlight important open research questions in this field. Throughout, we limit our scope to the benchmarking of gate-model quantum computers built from qubits (the primary quantum computing paradigm), e.g., we do not discuss benchmarks relevant only to quantum annealers or analog quantum simulators.
Quantum computer benchmarking
The field of quantum computer benchmarking evolved out of quantum tomography and error characterization [31, 32, 33]. These are techniques for reconstructing quantum states and learning detailed error models for qubits and quantum computers. The earliest “quantum computers” were physics experiments on one or two qubits, and physicists used characterization methods to learn about how their systems behaved [34, 35]. But tomography of $n$ qubits is exponentially expensive in $n$, so as $n$ grew it became necessary to summarize the performance of prototype quantum computers using a handful of metrics measured by benchmarks [17, 18].
A quantum computer benchmark (or just “benchmark” hereafter) is defined by (i) a set of computational tasks together with a procedure for performing those tasks on a quantum computer, and (ii) a method for computing one or more performance metrics from the (classical) data produced by running those tasks (Fig. 1a). This definition for a benchmark encompasses most uses of the term in quantum computer science, and it defines our scope. For example, benchmarks that directly measure materials properties (e.g., of an ion trap) are outside our scope.
For most benchmarks, the tasks comprise one or more quantum circuits, and for many benchmarks the output is a single, scalar metric of performance, e.g., quantum volume [36]. However, a single benchmark can compute multiple metrics (e.g., a success rate and a run time), or non-scalar quantities such as capability regions[37] (see Fig. 4). As with classical computers, no single benchmark or performance metric can capture all important aspects of a quantum computer’s performance.
Box 1. Errors in quantum computers. Gate-model quantum computers implement computations using quantum circuits, which are sequences of quantum logic gates [38] (see example below). When circuits are run on quantum computing hardware, errors can occur (red star, below). Errors cause quantum computations to fail by corrupting the outputs of quantum circuits, and most benchmarks quantify the magnitude and/or impact of these errors.[Uncaptioned image] A simple and popular model for errors in quantum circuits is that a random bit (and/or phase) flip occurs with some probability $\epsilon$ at each location in a circuit. So the probability of at least one error in a circuit of width $w$ and depth $d$ is $1-(1-\epsilon)^{wd}\lesssim wd\epsilon$, implying that a computation will likely succeed only if $\epsilon\ll 1/(wd)$. In this simple model, a quantum computer’s errors are described by a single error rate ($\epsilon$). But real quantum computers experience diverse and a priori unknown errors, and each kind of error has different effects [37, 39, 40]. Important examples of error types include coherent (i.e., systematic) and stochastic Pauli errors [39], crosstalk [41, 42], and slow drifts in a system’s parameters [43]. The complexities of the errors that occur in real quantum computers makes diverse, complementary benchmarks and metrics necessary.
From low-level to high-level benchmarks
The field of quantum computer benchmarking began in the early 2000s with the initial development of randomized benchmarking (RB) [17, 18, 19] as well as pioneering experiments on 2-12 qubits using quantum algorithms for solving tiny problems as informal benchmarks [44, 45, 46, 47]. The purpose of all these benchmarks was to directly quantify the rates of errors associated with qubits and gates, and their impact on quantum circuits (Box 1 reviews errors and circuits). For example, RB measures the error rates of quantum gates. These earliest methods are low-level benchmarks (see Fig. 2) that isolate qubit and gate performance from classical components (e.g., compilers) within an integrated quantum computing system. Since 2005, many low-level benchmarks that measure gate and circuit error rates have been (and continue to be) developed. They complement these earliest methods or address their limitations, and we discuss several of them in the next section.
In the last decade, benchmarking research has increasingly focused on high-level benchmarks (see Fig. 2) that quantify the overall performance of integrated (a.k.a. full-stack) quantum computers. The most well-known high-level benchmark is IBM’s quantum volume benchmark [36]. High-level benchmarks still predominantly quantify the impact of errors in quantum gates and qubits, but they do so by measuring their effect within the context of a complete quantum computing “stack” (see Fig. 3) that potentially includes a lot of classical computing subsystems. High-level benchmarks can quantify the impact of errors from the perspective of a quantum computer user executing high-level quantum programs. But to do so, they mix together many different aspects of performance [30, 48], including the performance of purely classical subsystems, that may not all be relevant to the pursuit of quantum utility.
The trend towards high-level benchmarks and metrics over the last decade has been driven by hopes for near-term quantum utility with “NISQ” computing [49, 50, 51] (see Box 2), by the emergence of commercial quantum computing systems with tightly integrated classical software stacks [30], and by the needs of prospective quantum computer users with limited quantum computing expertise. In the last few years, this trend has culminated with the development of many benchmarking suites centered around quantum algorithms and applications [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 6, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Although these trends are suggestive of a technology that is at or near utility, achieving utility seems like it will require major technological advances, notably ultra-reliable logical qubits (see Box 2). As long as quantum computing is in its infancy, high-level benchmarks—including those based on algorithms and applications—should be designed, used, and interpreted with caution.
The diverse roles and influences of benchmarks
In principle, benchmarks are passive tools for measuring performance metrics, used to localize devices in “performance metric space” (see Fig. 1b). But benchmarks also exert sociological forces. They influence decisions both small (e.g., gate calibrations) and large (e.g., funding allocations), and drive resources towards optimizing performance on those benchmarks. Benchmarks are used to compare competing quantum computing technologies and make judgements (sometimes implicit) about the current and future merits of each technology [53]. They therefore influence the design of future quantum computing systems. Benchmarks are also used to optimize individual quantum computing systems, by choosing system settings (e.g., compilation parameters, routing algorithms, or gate pulses) that maximize performance on a benchmark [48]. This optimization can even be an automated part of regular calibration or tune-up routines [73, 74].
Figure 2: Kinds of benchmark. Benchmarks vary widely in the abstraction level of their tasks, ranging from executing specific low-level quantum circuits to solving a computational problem, and by the complexity of the object whose performance they measure, ranging from individual logic gates to entire computing systems. The abstraction and complexity of a selection of important benchmarks and benchmark families are shown here.
Using benchmarks to inform a quantum computing system’s design and settings can easily lead to bad choices, because it is unlikely that any small set of benchmarks can summarize all important aspects of performance. Fundamental design decisions primarily aimed at maximizing performance on any existing high-level benchmarks would be particularly shortsighted, because no existing high-level benchmarks have been shown to reliably quantify progress towards quantum utility. For example, quantum volume favors systems with high qubit connectivity [30, 48] but higher connectivity will not necessarily enable better logical qubits. Optimizing an existing system’s settings to maximize its performance on benchmarks is typically less consequential, but can also result in poor performance unless the benchmarks are matched to the intended uses of that system. For example, using (only) one- and two-qubit RB to calibrate the gates of a system intended to run many-qubit circuits (e.g., NISQ algorithms or QEC) is a bad choice if that system has significant crosstalk errors that could be reduced by calibrations, because these benchmarks are largely insensitive to these errors [75, 24].
Widely adopted benchmarks and metrics that become de facto or official standards are particularly influential. Today, quantum volume is the most popular metric for summarizing an integrated quantum computer’s overall performance [76, 48]. A quantum computer’s total number of qubits and its average gate error rate (measured using RB) are the most popular metrics for summarizing its low-level performance. These three metrics are informative, but any small set of metrics has limitations. These metrics alone do not reliably quantify progress towards quantum utility (see later), implying that caution is necessary when using them to compare the quantum computing systems created by different companies and research groups.
The broad influence of benchmarks makes it critically important to design, adopt, and carefully use “good” benchmarks. Good benchmarks aid progress towards quantum utility (and eventually towards increasing utility) whereas bad benchmarks impede it. Benchmarks that measure technology-specific or highly technical metrics (e.g., leakage rates [77]) can and do aid this progress. But benchmarks that are technology independent and create easily digestible performance summaries (like quantum volume) will exert greater influence (for good or ill). Currently, the merits of those benchmarks should be primarily judged by whether they push the entire field of quantum computing towards computational utility.
Properties of good benchmarks
Good benchmarks are:
  1. 1.
    Well-motivated. A benchmark should measure well-motivated metrics of performance.
  2. 2.
    Well-defined. A benchmark should have an unambiguous procedure, i.e., any unspecified steps in that procedure should be intentional configurable parameters (e.g., some benchmarks allow creative compilation—see Box 4).
  3. 3.
    Implementation-robust. It should not be possible to exploit (“game”) a benchmark’s configurable parameters to obtain misleading results.
  4. 4.
    System-robust. A benchmark’s results should not be corrupted when the tested quantum computer experiences a priori unknown (small) errors.
  5. 5.
    Efficient. A benchmark should use a reasonable amount of all resources (e.g., quantum and classical computer time).
  6. 6.
    Technology independent. A benchmark should specialize to particular technologies or architectures only inasmuch as its metrics are only relevant in those contexts.
These properties are not binary, and not every useful benchmark scores highly on all of them. For example, informal benchmarks are often created to demonstrate experimental advances and these benchmarks need only serve the scientific purposes at hand (so, e.g., implementation robustness is irrelevant). However, for a benchmark to warrant trust from a broad cross-platform community, it needs to satisfy the above properties. Otherwise, that benchmark may mislead and contribute to misdirection of resources. Few (and perhaps none) of today’s high-level benchmarks satisfy all the criteria listed above.
Benchmarks must be robust if they are to reliably compare different technologies, or positively influence engineering decisions. Benchmarks are ideally designed to measure an independently defined (and well-motivated) metric (e.g., gate fidelity), and a robust benchmark accurately measures the intended metric under broad conditions. However, proving that a benchmark is robust is often difficult, and this is the central task in much of the theory of benchmarks. The accepted standard is to prove that a benchmark accurately measures its metric for any system whose qubits experience small Markovian errors [39], as exemplified by the theory of RB [28, 78, 79, 80]. However, real physical qubits (and likely logical qubits) often experience significant non-Markovian errors, so theories that address more general errors (and perhaps reveal weaknesses in existing benchmarks) are a need for the field.
Box 2. NISQ and fault-tolerant quantum computing. Quantum algorithms can be executed directly on physical qubits, which are not inherently fault tolerant. This approach, using systems akin to those available today, is referred to as noisy intermediate-scale quantum (NISQ) computing [49]. NISQ computers possess a moderate number of qubits, ranging from approximately 50 to 1000, and exhibit moderate error rates, between about $1\%$ and $0.1\%$. It is hoped that NISQ computers will demonstrate quantum utility through the use of heuristic algorithms specifically designed for such systems, commonly known as NISQ algorithms [49, 50, 51]. However, achieving utility-scale instances of quantum algorithms that exhibit a clear quantum advantage, such as Shor’s algorithm [81], is expected to require billions of gates [7, 8, 9, 10]. Consequently, executing these algorithms successfully necessitates error rates around $10^{-9}$ or lower, a target currently considered infeasible for physical qubits. Quantum computations can be rendered tolerant to errors by employing redundantly encoded logical qubits, protected using fault-tolerant quantum error correction (QEC) techniques [14]. According to fault tolerance theory, these logical qubits can achieve significantly lower error rates than their physical counterparts, provided the errors in the physical qubits are sufficiently rare and exhibit desirable characteristics (e.g., are adequately uncorrelated) [82]. In fault-tolerant quantum computing, physical qubits undergo repeated cycles of QEC, which differ significantly from the operations used in NISQ algorithms. This distinction necessitates different benchmarks and metrics for evaluating prototype NISQ and fault-tolerant quantum computers. Quantum algorithms are compiled very differently for NISQ and fault-tolerant architectures, and typically require many more gates (e.g., 50×50$\times$50 × more) for fault-tolerant execution. While most physical qubits can be directly manipulated using a continuous, universal set of one- and two-qubit gates [38], logical qubits are restricted to a discrete and non-universal set of “easy” gates, as per the Eastin-Knill theorem [15]. Achieving universal fault-tolerant quantum computation requires the use of “hard” gates, which require additional resources for implementation (e.g., magic state distillation [16]). This distinction has significant implications for the development of benchmarks for quantum computers.
Many high-level benchmarks measure self-defined metrics (e.g., quantum volume is defined as the quantity measured by its eponymous benchmark [36]) and this complicates assessing those benchmarks’ robustness. Such benchmarks are always intended to measure proxies for some independent (if perhaps imprecisely defined) aspect of performance. For example, high-level benchmarks based on one or more algorithms are typically meant to quantify how well a system can run “interesting” instances of those algorithms, but they do not necessarily do so. The intended interpretation of a benchmark implies an imprecise notion of robustness that can be assessed. Arguably, the quantum volume benchmark is robust [30, 48], whereas many other existing high-level benchmarks are not. However, the inherent imprecision and subjectivity in assessing the merits of benchmarks with self-defined metrics is unfortunate. The field would benefit from the creation of high-level benchmarks that robustly measure well-motivated and independent performance metrics.
Figure 3: How benchmarks interact with integrated quantum computers. Benchmarks test the joint performance of one or more parts of an integrated quantum computer’s “stack” (its qubits, compilers, routers, etc). They do so by inserting tasks into one level of the stack, and then analyzing output from the same (or a lower) level of the stack. Benchmarks can limit or adjust what each layer of the stack does (e.g., limiting the types of compilation), which can enable robust and efficient benchmarking. Benchmarks that enter and exit the stack at different levels measure fundamentally different aspects of performance, and form different categories of benchmark. Four important categories are shown here.
Benchmark efficiency is necessary to make a benchmark practical, but it is surprisingly challenging to design efficient benchmarks. Many benchmarks are intended to be run on $n$ qubits for any user-specified $n$, and such a benchmark is formally efficient (or scalable) if it requires resources that grow as a polynomial in $n$. But, in practice, a stronger but less precise notion of efficiency (scalability) is often desirable: the resources required are small for all relevant $n$. The creation of efficient benchmarks is an increasing focus of the benchmarking research community, due to the ongoing rapid increase in the number of qubits in quantum computers. Scalable benchmarks include modern versions of RB [22, 23, 24, 25], cycle benchmarking [83], and mirror circuit benchmarks [37, 84]. However, there are many widely-used benchmarks (e.g., cross-entropy benchmarking [XEB] [85] and the quantum volume benchmark [36]) that are not scalable, because they rely on classical simulations of general quantum circuits (see Box 3).
Kinds of quantum computer benchmark
There are many different kinds of quantum computer benchmark, and a bewildering array of specific benchmarks. We think that they can be best categorized and understood by how they interact with integrated quantum computers. Contemporary (prototype) quantum computing systems consist of multiple layers of hardware and software that is often called the quantum computing “stack” (see Fig. 3) in analogy to “software stacks” in classical computing [30]. How a benchmark interacts with this stack defines what parts of the quantum computer it tests, what performance metrics it measures, and whether the benchmark measures those metrics effectively.
Benchmarks and the quantum computing stack
Quantum computing systems can be divided into interacting subsystems in many ways, but in our context it is most useful to adopt a stack that reflects the flow of information though a computation (Fig. 3). Although different quantum computing systems can have very different architectures, most systems’ information flow is approximately as follows. A quantum computing system’s input is a computational problem (e.g., “factor 15”) and its output is classical data that is intended to solve this problem (e.g., $3\times 5=15$). The output is computed by running quantum programs on qubits, wrapped within some potentially complex classical computing systems (that may be automated or human-aided).
A quantum computer’s classical computing systems select an algorithm and then realize it as a quantum program. Here, a quantum program means a set of quantum circuits (see Box 2) that are perhaps wrapped inside a classical program, as in hybrid algorithms [49, 50, 51]. That quantum program is typically a high-level quantum program, i.e., it is written in a high-level language and its quantum circuits contain operations that are not native to that system’s computational qubits (which could be physical qubits or logical qubits built out of physical qubits running QEC). For example, it might use circuits that contain two-qubit gates between any pair of qubits (many systems have restricted connectivity), or large subroutines like the $n$-qubit quantum Fourier transform. This high-level program therefore typically needs to be compiled into a low-level quantum program whose circuits contain only gates that are “native” to that system’s computational qubits, before it can be executed. Finally, the system executes that low-level program, and processes the data. This conceptualization of the stack is an idealization, but it is a powerful aid for understanding and designing benchmarks.
Good benchmarks have a well-defined procedure and this requires that their interactions with the stack be precisely stated. Most existing benchmarks’ interactions with the stack fall within the following simple schema shown in Fig. 3. The benchmark inserts computational tasks into one layer of the stack (the input layer) and those tasks propagate down the stack until they reach the benchmark’s output layer. That output is then analyzed to compute performance metrics. For example, the quantum volume benchmark inserts high-level circuits (which the system compiles and runs) and it analyzes the resultant data to estimate the system’s quantum volume [36]. Some benchmarks also specify limitations on what each layer of the stack is allowed to do (benchmark rules in Fig. 3), and can even modify the intermediate outputs of the stack (output modifiers in Fig. 3). These esoteric aspects of a benchmark can be critical to its robustness or efficiency, e.g., output modifiers can circumvent the problem of efficiently verifying the outputs of general quantum circuits (see Box 3).
A benchmark quantifies the integrated performance of all the computational stages in between (and including) its input and output layers. Benchmarks that enter and exit the stack at different points quantify fundamentally different aspects of performance, and they constitute different and complementary categories of benchmark. Any entry and exit point in the stack defines a category of benchmarks. For example, compiler benchmarks [86, 87] isolate and test compilation algorithms by inserting high-level quantum circuits into the stack, and then analyzing the low-level circuits the compiler produces (using metrics such as the number of gates in those circuits). Three other categories of benchmark are also shown in Fig. 3, and are discussed below.
Benchmarking problem solving capabilities
A quantum computing system’s problem solving capabilities can be directly quantified using computational problem benchmarks that (i) challenge it to solve computational problems and then (ii) quantify performance in terms of metrics for the quality of (and/or time to) the solution. As illustrated in Fig. 3, these benchmarks enter the stack at the top level, and exit it at the lowest level, i.e., they test an entire computing system and compute a holistic performance metric. One reason such benchmarks are appealing is that they can be applied to any computing system, enabling comparison between radically different kinds of quantum computers as well as direct comparisons with classical computers.
Any computational problem can be used to define a computational problem benchmark, but it is particularly appealing to use problems that are believed to be intractable for existing classical computers yet efficiently solvable with quantum algorithms. An example is the task “factor RSA-2048” with the binary performance metric “were the correct factors found?”. Benchmarks like this will become increasingly important once quantum utility is first achieved, but they provide few insights into the performance of today’s quantum computers that appear to be far from solving important classically hard problems (e.g., we can confidently predict that all of today’s systems will fail to factor RSA-2048). Computational problem benchmarks based on classically intractable problems must also address the problem of how to quantify the quality of a purported solution to that problem. This result verification is often challenging (see Box 3).
Most of today’s algorithm-based benchmarks are not computational problem benchmarks [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 6, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], i.e., they do not challenge a system to simply solve a computational problem like “factor 15”. This is because today’s quantum computers cannot solve useful classically intractable problems, and any classically tractable problem can be solved by a quantum computer using primarily (or even only) its classical computing resources. This is illustrated by a series of high-profile experiments that factored 15 using a compilation of Shor’s algorithm that was only possible because 15’s factors were already known [88]. So, although quantum computing hardware capabilities are often demonstrated by running quantum algorithms on simplified problem instances [89, 90, 91], these experiments are rarely intended to define formal benchmarks. A good benchmark based on solving classically tractable problems must restrict the kinds of classical computation that a quantum computer is allowed to use when running the benchmark (see Box 4). For this reason, most algorithm-based benchmarks that are used today are high- or low-level program benchmarks, which are two categories of benchmark that we now discuss.
Box 3. The verification problem in benchmarking. Most benchmarks quantify how well a quantum computer executed its tasks, but designing benchmarks that measure such metrics efficiently is difficult in general. Computational problem benchmarks (see Fig. 3) need to quantify the accuracy or correctness of a purported solution to their problem(s). The correctness of the solutions to some classically intractable problems (those in the co-NP complexity class, like factoring) can be easily validated or falsified, but this is atypical of problems with known quantum speedups. In those cases, natural metrics for solution quality will not typically be efficient to measure, and so a good benchmark must instead measure an efficient-to-estimate and reliable proxy for those metrics. This was the approach taken in Google’s random circuit sampling (RCS) benchmark experiments [2]. Program and circuit benchmarks (see Fig. 3) face a similar verification challenge: many natural metrics for quantifying how well a quantum computer executed a quantum circuit compare the observed and error-free outcome distributions of that circuit [60, 84]. However, directly calculating these metrics requires classically computing the error-free outcome distribution, which is exponential expensive in the number of qubits in general. Many benchmarks avoid this problem by using circuits that are efficiently simulable classically [20, 22, 23, 24, 25]. However this approach is typically inappropriate for high-level program benchmarks, because efficiently simulable circuits are often particularly easy to compile into shallow (or even trivial) circuits (see Box 4). One solution to this challenge is to define high-level program benchmarks using circuits that are potentially intractable to simulate classically, and to then “intercept” the compiled low-level circuits before they are run (see output modifiers in Fig. 3), replacing them with proxies for those circuits whose performance is provably similar but that are efficiently simulable [84, 92, 93].
Benchmarking circuit execution capabilities
Quantum computations are implemented by running quantum programs, and so many benchmarks directly test a systems’ program-running capabilities. They typically do so by (i) tasking a quantum computer with running a set of quantum programs, and (ii) computing a metric that quantifies the error in those program’s execution (Box 3 discusses the challenge of efficiently computing such metrics). Because the core of any quantum program is the execution of quantum circuits, many of these benchmarks’ tasks are simply quantum circuits rather than more complex programs. Program (or circuit) benchmarks vary in the amount of compilation they allow (see Box 4), i.e., how high up they enter the stack, and they can be broadly categorized as high-level or low-level program benchmarks depending on whether they specify high-level programs that are then compiled by the system being tested, or low-level programs that permit little or no compilation.
High-level program benchmarks (see Fig. 3) permit broad optimizations (compiling, routing, etc) of their programs or circuits. The most widely-used such benchmark is the quantum volume benchmark, which quantifies a system’s performance on random circuits containing random two-qubit gates coupling arbitrary pairs of qubits [36]. In contrast, many of the recently-developed high-level program benchmarks are based on applications and algorithms and therefore use highly structured programs [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 6, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Prominent examples include the QED-C’s benchmarking suite [60, 52] and SuperMarQ [55]. These suites consist of an assortment of benchmarks each of which is based on an algorithm, including algorithms that are illustrative but unlikely to be useful (e.g., the Bernstein–Vazirani algorithm) [60], heuristic NISQ algorithm [55, 67], algorithms that are likely to require a fault-tolerant architecture to provide utility (e.g., Shor’s algorithm) [60], and algorithms that prepare canonical quantum states like GHZ states [55].
High-level program benchmarks are increasingly being used to compare different quantum computers [55], because they appear to enable simple and fair comparisons of systems with very different architectures. However, these comparisons can easily mislead. Contemporary quantum computers have not achieved quantum utility, so assessing these systems’ program-executing capabilities is useful only in so much as it assesses or incentivizes progress towards the goal of quantum utility. But improved performance on high-level program benchmarks can be obtained via system improvements that are unlikely to bring quantum utility closer. For example, many algorithms are unlikely to provide utility without fault tolerance, so an improvement to a quantum computing system’s classical compilation algorithm that enables better NISQ implementations of those algorithms (and therefore improves performance on existing algorithmic benchmarks) does not indicate progress towards quantum utility. The development of high-level program benchmarks that can reliably quantify, incentivize, and compare the progress of disparate architectures (e.g., NISQ versus fault-tolerant architectures) towards computational utility would be extremely valuable for the field.
Low-level program benchmarks (see Fig. 3) forbid intrusive classical compilation of their programs, and this prevents improvements on these benchmarks being driven by unimportant compilation algorithm optimizations. Examples of low-level program benchmarks include many RB methods [17, 18, 19, 20, 22, 23, 24, 25, 26, 27, 28, 83], XEB [85], mirror circuit benchmarks [24], and some algorithm-based benchmarks. Low-level program benchmarks typically permit at most localized compilation of their circuits, meaning replacing each individual gate in a circuit with sequences of native gates that synthesize that gate (this kind of limited compilation, discussed in Box 4, is often described in terms of compilation barriers between circuit layers). Low-level program benchmarks directly quantify the performance of a system’s qubits and gates, so they are they valuable for discovering or quantifying unaccounted-for errors in qubits and gates, by comparing actual performance with that predicted by a model [37, 75, 24]. Some low-level program benchmarks compute holistic performance metrics, such as capability regions [37] (see Fig. 4) that directly quantify circuit execution error, but many of them are designed to extract error rates for individual logic operations.
Box 4. The role of compilation in benchmarking. Quantum programs can be written at many levels of abstraction, and the level of abstraction used in a benchmark’s programs fundamentally impacts what it measures. The core of any quantum program is one or more quantum circuits, and a circuit is expressed using gates from some set (a.k.a. basis). This gate set could contain high-level gates (e.g., the $n$-qubit quantum Fourier transform), or only canonical one- and two-qubit gates (like CNOT and Hadamard gates), or only a specific system’s low-level gates (e.g., cross-resonance gates between pairs of connected qubits). Benchmarks often allow their circuits to be compiled into different gate sets, as this is necessary to make a benchmark executable on many different systems. A quantum program refers (often implicitly) to an equivalence class of programs, i.e., “run this program” means “run any program in this program’s family”, and a well-defined high- or low-level program benchmark specifies the equivalence class for each of its programs. Integrated quantum computers can contain powerful classical computers, and effective, robust benchmarks must take this into account when defining the kinds of compilation that are allowed. Many benchmarks are highly permissive, e.g., each of the quantum volume benchmark’s circuits $C$ can be replaced by any circuit $C\textquoteright$ ’ that (in the absence of errors) implements (approximately) the same unitary as $C$. This is a common approach, as benchmarks like this jointly test a system’s compilation algorithms and qubits. A similar and conceptually simpler approach is for a benchmark to allow its circuits to be replaced with any other circuits that (in the absence of error) produce samples from the same probability distribution, but such benchmarks can be gamed. This is because (i) the benchmark permits replacing each of its circuits $C$ with a potentially trivial circuit $C^{\prime}$ that simply encodes the already-computed output distribution of $C$, and (ii) this is feasible to do (i.e., a system’s classical compilers could find such a $C^{\prime}$) whenever each $C$ can be quickly simulated classically.
Benchmarking components and subroutines
Many of the most mature benchmarks are designed to measure the error rates of the fundamental logic operations from which quantum circuits are built (e.g., individual single-qubit and two-qubit gates, layers of gates, measurements, etc). These component benchmarks (see Fig. 2) typically achieve this by running low-level circuits containing the components of interest and inferring how those components performed from the performance of those circuits. The paradigmatic component benchmarks are the RB protocols [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
RB runs low-level circuits containing random gates, which average out the details of those gate’s error processes. This causes the success rate of these circuits to decay exponentially in circuit depth, and fitting this data to an exponential enables estimating those gates’ mean error rate. The de facto standard RB method measures the mean error rate of the set of all one- or two-qubit Clifford gates [20], and there is now a large family of RB methods and closely-related techniques, which adapt or improve standard RB in various ways. Prominent examples include XEB [85], direct RB [21], mirror RB [24, 25], binary RB [22], interleaved RB [29], cycle benchmarking [83], and character RB [27, 28]. XEB, direct RB, mirror RB, and binary RB are closely-related methods for measuring the average error rate of a set of layers of native gates. Interleaved RB and cycle benchmarking are methods for estimating the error rate of a single gate or layer of gates. Character RB enables adapting standard RB to sets of gates that form groups other than the Clifford group.
Benchmarks that can measure the performance of fundamental operations on logical qubits will soon become increasingly important. Many of today’s benchmarks can be applied to logical qubits [26], but most of them were primarily designed and analyzed for benchmarking physical qubits. The degree to which the theory and assumptions underpinning existing benchmarks (e.g., approximate Markovianity [39]) will apply to logical qubits is currently unknown. However, we anticipate that new component benchmarks that are better suited to the properties of logical qubits will be needed. For example, some operations that logical qubits will perform are fundamentally different to any operations performed on physical qubits (e.g., lattice surgery [94]) and benchmarks that measure the performance of these complex computational primitives will be needed.
The move towards fault-tolerant architectures also brings immediate needs for component benchmarks that focus on the properties of physical qubits and gates that are of most relevance in this setting. Many benchmarks measure the total rate of errors in fundamental logic operations (e.g., layers of gates), which is predictive of the failure rates of circuits executed directly on physical qubits and is therefore particularly relevant in a NISQ setting. However, some kinds of errors are much more damaging than others in fault-tolerant architectures, because they are more costly to correct. Similarly, some fundamental logical operations and subroutines are central to fault-tolerant quantum computing but of little or no importance in a NISQ architecture. Important examples include mid-circuit measurements, parity checks, and syndrome extraction cycles. Benchmarks for these primitives are relatively underdeveloped. Recent breakthroughs demonstrating components of fault-tolerant quantum computing [3, 95, 96, 97] introduced some prototype benchmarks, but mature benchmarks for the physical primitives of fault-tolerant quantum computing are a near-term need for the field.
Measuring progress to quantum utility
We think the most compelling current purpose of quantum computer benchmarks is measuring progress toward quantum utility. Correctly designed and interpreted benchmarks can help stakeholders understand and quantify the value of steps toward that goal. However, “quantum utility” is not a uniquely defined goal. There are many important, as-yet-unsolved problems that a quantum computer might solve, and many possible roadmaps to creating such a computer. We expect tracking progress to require a range of complementary benchmarks that are motivated by challenge problems, and interpreted relative to resource estimates and hardware roadmaps.
A challenge problem is a specific instance of an important computational problem that, if solved by a quantum computer, would constitute quantum utility. Challenge problems turn “quantum utility” into concrete computational goals. A resource estimate is a precise accounting of the computational resources required to solve a challenge problem. Resource estimates map computational goals to engineering goals. A hardware roadmap comprises plans and schedules for a sequence of increasingly capable quantum computers culminating in one that can solve a challenge problem. Hardware roadmaps define paths along which progress is tracked.
These ingredients enable choosing, creating, or adapting specific benchmarks that can test whether each milestone on a roadmap has been achieved, measure progress between milestones, and (perhaps) forecast the cost of future progress. In their absence, benchmarking cannot effectively track progress. Benchmark results can show that some property has improved, or that one device exceeds another in some way, but they can’t quantify the value of that change.
Challenge problems for quantum computing
Quantum computers will not outperform classical computers for all computational tasks, but there’s good reason to believe they can speed up specific structured tasks. If such a task is also (subjectively) useful, it can represent quantum utility. A challenge problem is a specific instance of such a task, crafted to be (1) feasible with a quantum computer, (2) infeasible—or at least more costly—with any other computer, and (3) useful. Plausible challenge problems are sufficiently rare and distinct that it makes sense to treat each one individually.
The best candidate challenge problems involve calculating quantities that can’t feasibly be computed without a quantum computer. Although problems that are just very hard for classical computers and much easier with quantum computers are also appealing, they are moving targets. Because classical algorithms can often be improved, and the power and speed of classical computers grows steadily over time, even a 1000×1000$\times$1000 × quantum advantage could be erased within years (or even days, with rapid classical algorithm development to counter a claimed quantum advantage [98, 99, 100]). So the most compelling challenge problems are specific instances of computational problems for which the quantum computer has exponential advantage in a key resource (generally time-to-solution).
Two of the best studied categories of challenge problems are integer factorization and quantum chemistry. Instances of factorization are easy to describe (e.g., “factor this 2048-bit semiprime” [7]), they scale naturally in size and difficulty, and their solutions can be verified easily and rigorously. Instances of quantum chemistry can be stated precisely (e.g., “sample from the energy eigenspectrum of the FeMo cofactor to at least chemical accuracy” [8]), but their solutions are typically only heuristically verifiable, and “hardness” is challenging to quantify systematically even though a domain expert can usually assess a given instance’s difficulty.
Few sharply defined challenge problems like factoring and chemistry are known. Establishing more would be very useful, but appears to be very hard. This has fostered confusion about how soon quantum utility might be achieved. The few known challenge problems suggest that quantum utility is very difficult, and will not be achieved soon. But there is a huge gray area containing important problems that might admit quantum speedups and quantum-friendly problems that might be useful. Identifying even one additional unambiguous challenge problem from among them could change the benchmarking landscape drastically.
Discrete optimization, machine learning, and linear algebra offer promising candidate challenge problems, but it’s not clear whether useful speedups even exist in these areas. If they do, the resulting challenge problems will probably resemble quantum chemistry (difficult to systematize) more than factoring (sharp and verifiable). Factoring may be unique among known challenge problems in the concision and clarity of any stated instance.
To get more good challenge problems, we may need to rely on heuristic notions of verification and difficulty. If so, consensus will only emerge from detailed analyses of the best classical algorithms’ capabilities and substantive debate between domain experts. Recent developments in quantum chemistry illustrate this process. Active debate about what instances are “hard” [101], and about which quantum advantages will be both feasible and useful [102], has proceeded in parallel with research advances in simulation algorithms and careful constant-factor resource estimates [103, 8, 104] that are necessary to define and assess quantum utility.
Figure 4: Assessing quantum computer performance via capability. This figure illustrates one way to compare experimentally benchmarked performance against resource estimates for challenge problems, using a multidimensional capability metric. Challenge problems and benchmark tasks are represented by the width (some measure of the number of qubits) and depth (some measure of the number of clock cycles) of a quantum circuit that performs the task. Regions indicate the circuits performable by two real-world quantum computers—Google’s Sycamore (green) as extrapolated from results in Arute et al. [2], and an ensemble of IBM Q devices (pink) benchmarked by our group [37]—and one hypothetical quantum computer (blue) (we use a success threshold of $1/e$). Points indicate constant-factor resource estimates for three candidate challenge problems analyzed in the literature [7, 8, 9]. For these problems, width is the number of logical qubits, not accounting for logical qubits used in distillation or routing, and depth is the total number of non-Clifford operations (i.e., Toffoli and/or T gates). These metrics are somewhat crude, but indicate the rough scale of resources required for these challenge problems. We emphasize the wide gulf between that “utility” scale and current state of the art capabilities—logarithmic axes were required to compress both scales into one figure. Plots like this one could enable stakeholders to track and extrapolate the growth of quantum computer capabilities over time, toward eventual achievement of quantum utility.
Resource estimates for challenge problems
Resource estimation addresses the question “How big, fast, and reliable would a quantum computer need to be?” in order to solve a specific challenge problem [103, 105, 106, 107, 7, 8, 108, 109, 104, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 9]. A resource estimate for a challenge problem is a description of a minimal (known) quantum computer that would suffice to solve the challenge problem efficiently. We say “minimal” rather than “minimum” because some resources can be traded for others (e.g., memory vs speed, or space vs time).
Quantum algorithms are typically analyzed in terms of their asymptotic scaling, e.g., the manner in which the requisite quantum computational resources scale with problem size, solution accuracy, etc. Resource estimation often carries the connotation that the constant prefactors for specific problem instances are also calculated. Resource estimates are quantitative, and indicate whether theoretical quantum speedups are practically feasible. For example, recent work [131, 132] provides a strong argument that good challenge problems are unlikely to involve merely quadratic quantum speedups, even under optimistic assumptions about advances in quantum hardware and architectures and relatively pessimistic assumptions about the performance of classical hardware.
Resource estimates can be specified at various levels of abstraction and detail, and different levels of detail can motivate entirely different benchmarking paradigms. The purpose of a resource estimate, at any level of detail, is to describe the key properties and performance characteristics of a plausible quantum computer that could solve a challenge problem.
A simple logical resource estimate might be “$N$ qubits, capable of executing [specific gates], with error probability $\leq\epsilon$ per gate, running for time $T$,” derived directly from a fully-compiled algorithm known to solve the problem. Logical resource estimates for most known challenge problems mandate error probabilities of $\epsilon<10^{-15}$/gate, and error rates this low are widely believed to be achievable only via fault-tolerant QEC.
More granular resource estimates account for the unavoidable existence of “hard” and “easy” logic operations in every fault-tolerant architecture (see Box 2). They count the minimum number and cost of “hard” operations that will need to be implemented using tricks like magic-state distillation and injection [16] or code switching [133]. A more detailed approach specifies a particular QEC code and method of implementing fault-tolerant computation within that code (e.g., lattice surgery [94]).
The most sophisticated physical resource estimates provide a full architectural specification—“$N^{\prime}$ physical qubits, capable of implementing [specific fault-tolerant quantum computing scheme], with error probability $\leq\epsilon^{\prime}$ per physical gate, running for time $T^{\prime}$.” Today’s best physical resource estimates use oversimplified error models, and future resource estimates may treat the physical hardware in even greater detail in order to specify engineering specs like materials quality or temperature.
Roadmaps to solving challenge problems
Resource estimates (for specific challenge problems) tell us where the goalposts of quantum utility are. Roadmaps define the distance to those goalposts. Tracking progress toward a specific computationally useful computer requires knowing what progress looks like. A roadmap describes a sequence of increasingly capable prototype devices. Some steps in this sequence may introduce key technologies that don’t immediately increase computational power. Roadmaps enable benchmarking to recognize such steps as progress and quantify their value.
Each prototype can be associated with a list of measurable specifications, e.g., gate error rates, latency times, crosstalk rates, control bandwidth, etc. A roadmap is (at minimum) a sequence of such specification lists, which trace out a path to a useful device. Real devices are enormously complex, with many specifications. To track progress along a roadmap using benchmarks, a short list of key specifications needs to be identified. Those specifications should be (i) measurable or inferrable using benchmarks, and (ii) sufficient to prove that the prototype achieves its goals within the context of the roadmap.
The most relevant specifications (and thus the benchmarks necessary to infer them) will change at different points along a roadmap. For example, in the early stages of a roadmap toward a fault-tolerant quantum computer with millions of qubits [7, 8, 9], gate error rates on physical qubits may be paramount and require direct benchmarking. In later stages, after multiple fault-tolerant logical qubits have been assembled, logical qubit error rates may replace them as a key specification.
Benchmarking progress towards utility
In the context provided by a challenge problem, a resource estimate, and a roadmap, benchmarks can be constructed for the purpose of measuring the specifications of a physical quantum computer (i.e., a prototype), and identifying its position along the roadmap (see Fig. 1b). The best benchmarks for this task will vary with the scale of the prototype and the details of the roadmap.
One way to measure a prototype’s performance relative to a roadmap is to use benchmarks that isolate properties of specific low-level components (e.g., error rates of physical one- and two-qubit gates) and combine them using calculations or simulations to infer the values of higher-level metrics (which can then be compared to target values) such as logical qubit error rates. This approach is implicit in fault-tolerance threshold theorems that compute high-level computational properties from low-level models. However, quantum computers are complex systems, whose emergent behavior may not be reliably captured or predicted by modeling informed by component benchmarks [37, 25, 75, 24]. Holistic benchmarks (see Fig. 2) offer a complementary approach. They can probe computational power directly, provide direct evidence of progress along a roadmap, and can be used to calibrate simulations.
It is unlikely that any single scalar metric will accurately assess or incentivize progress toward quantum utility. However, holistic benchmarks can report rich multi-dimensional metrics with greater descriptive power. A particular example that we have found useful and inspiring is capability benchmarks [37]. Since a quantum computer’s computational power comes from running programs, one way to probe that power is to ask “What programs can it run?” The set of programs that a computer can run, with reasonable accuracy using reasonable time and energy, is called its capability [37, 134]. Formally, this set-valued metric is infeasible to measure or even write down because there are far too many programs. But it can be sketched or approximated using a short list of key program features. If sufficiently faithful, a sketch of capability provides intuitive heuristic answers to many questions about a quantum processor’s computational ability.
Capabilities can be sketched by choosing a limited set of program features, such as circuit width and circuit depth [135, 37] or the number of “hard” gates (e.g., two-qubit gates) in a circuit [6]. We define a circuit class containing all programs with the same values of the selected features. Ideally, we seek to choose features so that a given processor will be able to successfully run every circuit in a class, or none of them. Inasmuch as this holds, a benchmark of moderate complexity can probe the processor’s ability to run sample programs from each circuit class. The result is a high-dimensional metric that can be visualized as a capability region [37] (Fig. 4) in the space defined by the program features—and compared directly to resource estimates, for challenge problems, displayed on the same plot.
This specific approach has clear limitations that have not yet been overcome (What kind of programs should be run? Do good features exist? How many are needed? What defines “the same” program on different architectures? Should the programs be defined at high or low levels of abstraction, and what kinds of compilation should be allowed?). Perhaps an entirely different approach will work better. But we suspect that any successful strategy for benchmarking progress to utility will need to move beyond the single-scalar-metric paradigm, and will need to address these challenges or similar ones.
An open problem for any such approach is how to find or construct genuinely representative proxy programs that behave like programs that would solve a challenge problem, but that can be scaled seamlessly to fit on any given prototype. A related open question is how these programs change when adapted to different architectures, and how to construct benchmarks that evaluate both NISQ and fault-tolerant architectures on equal footing. We foresee a significant era during which advanced NISQ devices compete with, and should be fairly compared to, early fault-tolerant devices. They may pursue very different challenge problems via very different roadmaps. Creating benchmarks that can compare progress of all architectures in a shared context seems challenging, but valuable to many stakeholders.
Acknowledgements
This material was funded in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Quantum Testbed Pathfinder Program. T.P. acknowledges support from an Office of Advanced Scientific Computing Research Early Career Award. A.D.B. acknowledges support from the National Nuclear Security Administration’s Advanced Simulation and Computing Program and the Department of Energy (DOE) Office of Fusion Energy Sciences “Foundations for quantum simulation of warm dense matter” project. Sandia National Laboratories is a multi-program laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC., a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. All statements of fact, opinion or conclusions contained herein are those of the authors and should not be construed as representing the official views or policies of the U.S. Department of Energy, or the U.S. Government.
Author contributions
All authors contributed to developing the perspective presented here. TP and RBK led the writing of the manuscript, with all authors contributing.
Competing interests
The authors declare no competing interests.